Lee Holmes deposited $17,500 in a new savings account at 7% interest compounded semiannually. At the beginning of year 4, Lee deposits an additional $42,500 at 7% interest compounded semiannually.

At the end of 6 years, what is the balance in Lee’s account? (Use the Table provided.)
Answer #1

Annual interest of 7% means 3.5% of Semiannual interest or Effectivr interest rate is 3.5% for 6 months.

Period

Deposit balance

Addition made

Interest earned

Ending balance

Year 1

Half year

$ 17,500.00

$     612.50

$ 18,112.50

Half year

$ 18,112.50

$     633.94

$ 18,746.44

Year 2

Half year

$ 18,746.44

$     656.13

$ 19,402.56

Half year

$ 19,402.56

$     679.09

$ 20,081.65

Year 3

Half year

$ 20,081.65

$     702.86

$ 20,784.51

Half year

$ 20,784.51

$     727.46

$ 21,511.97

Year 4

Half year

$ 21,511.97

$ 42,500.00

$ 2,240.42

$ 66,252.39

Half year

$ 66,252.39

$ 2,318.83

$ 68,571.22

Year 5

Half year

$ 68,571.22

$ 2,399.99

$ 70,971.21

Half year

$ 70,971.21

$ 2,483.99

$ 73,455.21

Year 6

Half year

$ 73,455.21

$ 2,570.93

$ 76,026.14

Half year

$ 76,026.14

$ 2,660.91

$ 78,687.05

Interest for year 4 first half year is on $21511.97+42500

Balance of deposit at the end of Year 6 =$78,687.05

Alternate solution using Future value table.

Future value table is at interest rate of 3.5%

Deposit balance at the end of year 6

A

Initial deposit

$ 17,500.00

B

PV factor total at year 3

0.8135

C=A/B

Future value factor for 3 years

$ 21,511.99

D

Add: Additional deposit

$ 42,500.00

E=D+C

Total beginning balance of year 3

$ 64,011.99

F

Future value factor for 3 years

0.8135

G=E/F

Deposit balance at the end of year 6

$ 78,687.14

Balance of deposit at the end of Year 6 =$78,687.14

Answer by both alternative are different due round offs.

Please leave a comment if you need more clarification.

If answer does not match, just let me know so that I can help.